Let X1,X2,.....,Xn be random sample of size n from a distribution with probability 

density function 

f(X; ,0)={ (θX) to the power (θ-1) ,0<X<1 , θ>0 Obtain a maximum likeyhood

0, elsewhere.

estimator of θ.


1
Expert's answer
2021-06-07T14:53:51-0400

"f(X,\theta)=\theta X^{\theta-1}, 0\le X\le1, \theta>0"

"L(\theta)=\prod_{i=1}^nf(X,\theta)"

"=\theta ^n\prod_1^n(Xi)^{\theta-1}"

"\ln L(\theta)=n\ln\theta+(\theta-1)\ln \prod_1^n Xi"

To maximize "\ln L(\theta)", we equate the first derivative of "\ln L(\theta)"to zero and solve for "\theta"

"\frac{d}{d\theta}\ln L(\theta)=\frac{n}{\theta}+\ln \prod_1^n Xi"

"\frac{n}{\theta}+\ln \prod_1^n Xi=0"

"\frac{n}{\theta}=-\ln \prod_1^n Xi"

"\hat{\theta}=-\frac{n}{\prod_1^nXi}"